The Variables x and y vary inversely. y=3/5 when x=7/2 Find an equation that relates the variables
Answer by stanbon(48558)

The Variables x and y vary inversely. y=3/5 when x=7/2 Find an equation that relates the variables
---------------
y = k/x
(3/5) = k/(7/2)
k = (35/6)
------------------
EQUATION:
y = (35/6)/x
y = 35/(6x)
================
Cheers,
Stan H.